OpenStudy (anonymous):

Prove that P(n) : n(n+1)(n+5) is divisible by 3

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

What are our tool? Are you trying to do this with induction? or do we have modular arithmetic at our disposal?

OpenStudy (anonymous):

induction

OpenStudy (beginnersmind):

My favourite would be to write it as n(n+1)(n+2) + 3n(n+1) The second term is divisible by 3 and the first one is the product of 3 consecutive integers therefore also divisible by 3. The sum of 2 numbers both divisible by 3 is also divisible by 3.

OpenStudy (anonymous):

@beginnersmind i didn't get it :/ sorry can u explain it step by step ?

OpenStudy (anonymous):

According to induction: Put n = 1 and see are you getting what the question says..

OpenStudy (anonymous):

@beginnersmind do not stop. Carry on.. I am just trying as @ashna is doing..

OpenStudy (anonymous):

i know that water tell me from , to prove p(k+1) is true

OpenStudy (beginnersmind):

My answer didn't use induction so I'd rather not go into a long explanation.

OpenStudy (anonymous):

Ha ha ha ha... You knew that?? Just kidding..

OpenStudy (anonymous):

we get k = 3M/ (K+1)(K+5) Right ?

OpenStudy (anonymous):

Replace n by k+1 first...

OpenStudy (anonymous):

yeah did , then ?

OpenStudy (anonymous):

Then Look carefully it will also be divisible by 3.. Ha ha ha ha...

OpenStudy (anonymous):

c'mon Water i don't understand :I

OpenStudy (anonymous):

\[= (k+1)(k+2) (k+6)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Really??

OpenStudy (anonymous):

where r yu goin to substitute k = 3M/ (K+1)(K+5) ?

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

It is now 6 when I studied Induction..

OpenStudy (anonymous):

*6 years..

OpenStudy (beginnersmind):

Ok, this is how you do it with induction. First prove it for n =1 (plug it in and check if it's divisible by 3) Second assume that it's true for P(k). Using this try to prove it's also true for P(k+1) In this case I would try to prove that P(k+1) - P(k) is divisible by 3.

OpenStudy (anonymous):

yeah .. on assuming i got k = 3M/ (K+1)(K+5) 3rd step am stuck :I

OpenStudy (beginnersmind):

what does the M stand for?

OpenStudy (beginnersmind):

Ah, ok, see what you did there. You said there's a number M such that P(k) = 3M

OpenStudy (anonymous):

M = divisible by 3

OpenStudy (beginnersmind):

Ok, I'd do it slightly differently. I'd prove that the difference of P(k+1) and P(k) is divisible by 3. Then using this and the induction hypothesis it follows that P(k+1) is also divisible by 3. Does that make sense?

OpenStudy (anonymous):

yes :)

OpenStudy (beginnersmind):

Cool :) To check, what did you get for P(k+1) - P(k) ?

OpenStudy (anonymous):

What if we find the value of k+1 from the assumption??

OpenStudy (anonymous):

\[k+1 = \frac{3M}{k (k+5)}\]

OpenStudy (anonymous):